OpenStudy (anonymous):

((-16-(-2+1))x2)/-5 Need help on solving this don't just want the answer want to know how you did it. Thanks!<33

OpenStudy (lalaly):

1) -2+1=-1 2) -16-(-1) = -16+1 = -15 3) -15 x 2 = -30 4) -30/-5= 6

OpenStudy (precal):

Use order of operation. Do what is inside the parenthsis first (that is step one from lalaly)

OpenStudy (anonymous):

Oh okay thank you!! (:

OpenStudy (lalaly):

PEMDAS parentheses| exponents | multiplication & division from left to right | addition & subtraction from left to right

OpenStudy (anonymous):

got it<3

OpenStudy (lalaly):

;):D